OpenStudy (asapbleh):

Find all six trigonometric functions of theta if the given point is on the terminal side of theta. (3,4)

OpenStudy (asapbleh):

is this correct? \[\sin \theta=4/5\] \[\cos \theta=3/5\] \[\tan \theta=4/3\] \[\cot \theta=3/4\] \[\sec \theta=5/3\] \[\csc \theta=5/4\]

OpenStudy (anonymous):

Looks right to me.

OpenStudy (anonymous):